# Mitigating Graph Covariate Shift via Score-based Out-of-distribution Augmentation

### v1

@@ 5.3 OOD controlled graph generation
Table 3: Visualizations of the augmented GOOD-Motif-base graphs generated by OODA. The graphs $\tilde{G}$ with $\lambda=0.1$, $\lambda=0.2$, and $\lambda=0.3$ represent the OOD graphs generated by OODA under different values of $\lambda$.
6 Conclusion
In this work, we proposed OODA, an out-of-distribution graph generation framework based on a score-based diffusion probabilistic model, designed to address covariate shifts in graph learning. Our approach generates OOD graph samples that integrate both explored environmental and stable features, eliminating the need to separate them during training. Furthermore, OODA can simultaneously explore new environments in graph structure, node features, and edge features. While score-based diffusion models demonstrate significant potential in handling diverse covariate shifts, they present scalability challenges when applied to large-scale graphs. Additionally, generating OOD graphs may require careful tuning of hyperparameters in the guidance scheme to balance exploration quality, particularly across different datasets. In this study, we applied minimal hyperparameter tuning to achieve competitive results. Future work could focus on developing scalable diffusion models and exploring parameter-efficient tuning strategies to further enhance OOD graph generation.
Appendix A Appendix
A.1 Proofs
$\nabla_{G_{t}}\log p_{t}\left(G_{t},\mathbf{y}_{G}\mid\mathbf{y}_{\text{ood }}=\lambda\right)=\nabla_{G_{t}}\log p_{t}\left(G_{t}\mid\mathbf{y}_{G},\mathbf{y}_{\text{ood }}=\lambda\right)$
@@ A.1 Proofs
$\log p_{t}\left(G_{t},\mathbf{y}_{G}\mid\mathbf{y}_{\text{ood }}=\lambda\right)=\log p_{t}\left(G_{t},\mathbf{y}_{G},\mathbf{y}_{\text{ood }}=\lambda\right)-\log p_{t}\left(\mathbf{y}_{\text{ood }}=\lambda\right)$
Since $p_{t}\left(\mathbf{y}_{\text{ood }}=\lambda\right)$ is independent of $G_{t}$, $\nabla_{G_{t}}\log p_{t}\left(\mathbf{y}_{\text{ood }}=\lambda\right)=0$. Therefore,
$\nabla_{G_{t}}\log p_{t}\left(G_{t},\mathbf{y}_{G}\mid\mathbf{y}_{\text{ood }}=\lambda\right)=\nabla_{G_{t}}\log p_{t}\left(G_{t},\mathbf{y}_{G},\mathbf{y}_{\text{ood }}=\lambda\right)$
@@
$\log p_{t}\left(G_{t}\mid\mathbf{y}_{G},\mathbf{y}_{\text{ood }}=\lambda\right)=\log p_{t}\left(G_{t},\mathbf{y}_{G},\mathbf{y}_{\text{ood }}=\lambda\right)-\log p_{t}\left(\mathbf{y}_{G},\mathbf{y}_{\text{ood }}=\lambda\right)$
Since $p_{t}\left(\mathbf{y}_{G},\mathbf{y}_{\text{ood }}=\lambda\right)$ is independent of $G_{t}$, $\nabla_{G_{t}}\log p_{t}\left(\mathbf{y}_{G},\mathbf{y}_{\text{ood }}=\lambda\right)=0$. Therefore,
$\nabla_{G_{t}}\log p_{t}\left(G_{t}\mid\mathbf{y}_{G},\mathbf{y}_{\text{ood }}=\lambda\right)=\nabla_{G_{t}}\log p_{t}\left(G_{t},\mathbf{y}_{G},\mathbf{y}_{\text{ood }}=\lambda\right)$
@@
$\nabla_{G_{t}}\log p_{t}\left(G_{t},\mathbf{y}_{G}\mid\mathbf{y}_{\text{ood }}=\lambda\right)=\nabla_{G_{t}}\log p_{t}\left(G_{t}\mid\mathbf{y}_{G},\mathbf{y}_{\text{ood }}=\lambda\right)$
A.2 Graph covariate shift
In this work, we address the challenge of OOD graph classification, where the goal is to develop models trained on a dataset $\mathcal{D}_{\mathrm{tr}}$ that can generalize effectively to a dataset $\mathcal{D}_{\mathrm{te}}$. Considering invariant perspectives under covariate shifts (Gui et al., 2022; Sui et al., 2024), we note while $P_{\mathrm{tr}}(Y\mid G)=P_{\mathrm{te}}(Y\mid G)$ holds, the marginal distributions of graphs differ, i.e., $P_{\mathrm{tr}}(G)\neq P_{\mathrm{te}}(G)$. This discrepancy may stem from the training set’s limited size or diversity and the unpredictable conditions in test environments. For instance, acquiring sufficient samples of molecules for property prediction can be costly and challenging. Moreover, geometric deep learning models are frequently applied to predict properties of molecules with unseen scaffolds. Prior study (Li et al., 2024) identifies two primary types of covariate distribution discrepancies: (1) $P_{\mathrm{tr}}(\bm{X})\neq P_{\mathrm{te}}(\bm{X})$ while $P_{\mathrm{tr}}(\bm{A})=P_{\mathrm{te}}(\bm{A})$ and $P_{\mathrm{tr}}(\bm{E})=P_{\mathrm{te}}(\bm{E})$, exemplified by the GOOD-CMNIST dataset (Gui et al., 2022), where digits of different colors indicate different environments that correspond to dataset splits. (2) $P_{\mathrm{tr}}(\bm{A},\bm{E})\neq P_{\mathrm{te}}(\bm{A},\bm{E})$ or $P_{\mathrm{tr}}(\bm{A},\bm{X},\bm{E})\neq P_{\mathrm{te}}(\bm{A},\bm{X},\bm{E})$, as seen in the GOOD-Motif dataset (Gui et al., 2022) where training and testing environments differ in graph size and bases. In the GOOD-HIV benchmark (Gui et al., 2022), the molecular scaffolds differ between the training and testing datasets.
The covariate shift, whether in graph structure, node features or eddge features, poses significant challenges to OOD generalization. This work aims to provide a general framework to address all three types of covariate shifts.
A.3 Metrics for measuring distributional differences
In this section, we provide detailed implementation steps for measuring the distributional differences between the augmented dataset and the training dataset. Following (Thompson et al., 2022), we first use an untrained random GIN, $h$, to extract graph embeddings from both the augmentation distribution and the training distribution. The maximum mean discrepancy (MMD) is then computed to quantify the dissimilarity between the graph embedding distributions:
$\operatorname{MMD}^{2}(P\|\tilde{P})=\mathbb{E}_{g,\tilde{g}\sim P}[k(h(g),h(\tilde{g}))]+\mathbb{E}_{g,\tilde{g}\sim\tilde{P}}[k(h(g),h(\tilde{g}))]-2\mathbb{E}_{g\sim P,\tilde{g}\sim\tilde{P}}[k(h(g),h(\tilde{g}))]$
where $k(\cdot,\cdot)$ is the RBF kernel proposed by (You et al., 2018). As recommended by (Thompson et al., 2022), the MMD RBF scalar is also one of the most reliable metrics for measuring distributional differences:
@@ A.3 Metrics for measuring distributional differences
Additionally, we employ the Earth Mover’s Distance (EMD) from (Thompson et al., 2022) to compute pairwise distances $d(\cdot,\cdot)$.
A.4 Experimental Details
A.4.1 Dataset Details
We utilize six datasets from the GOOD benchmark (Gui et al., 2022), including GOOD-Motif-base, GOOD-Motif-size, GOOD-CMNIST-color, GOOD-HIV-scaffold, GOOD-HIV-size, and GOOD-SST2-length. The GOOD benchmark (Gui et al., 2022) is the state-of-the-art framework for systematically evaluating graph OOD generalization. It carefully designs data environments to induce reliable and valid distribution shifts. The selected datasets span a diverse range of domains, covering covariate shifts in general graphs, image-transformed graphs, molecular graphs, and natural language sentiment analysis graphs. The dataset details are as follows:
  • GOOD-Motif: GOOD-Motif is a synthetic dataset from Spurious-Motif (Wu et al., 2022) specifically designed to investigate structure shifts. Each graph consists of an environmental base graph connected to a label-determining motif. The two primary covariate shift domains are the base graph type and graph size. For base covariate shift, the training distribution includes graphs with wheel, tree, and ladder base structures, while the validation set features star base graphs, and the test set contains path base graphs. For size covariate shift, the training distribution consists of graphs with sizes ranging from $6$ to $45$ nodes, the validation set contains graphs with sizes between $20$ and $75$ nodes, and the test set comprises graphs with sizes ranging from $68$ to $155$ nodes.
  • GOOD-CMNIST: GOOD-CMNIST is a semi-synthetic dataset designed to investigate node feature shifts. It consists of graphs transformed from MNIST handwritten digit images using superpixel techniques (Monti et al., 2017). Node color features are manually applied, making the color shift environment independent of the underlying structure. Specifically, for covariate shift, digits are colored using seven different colors. The training distribution includes digits colored with the first five colors, while the validation and test distributions contain digits with the remaining two colors, respectively.
  • GOOD-HIV: GOOD-HIV is a small-scale, real-world molecular dataset sourced from MoleculeNet (Wu et al., 2018). The nodes in these molecular graphs represent atoms, and the edges represent chemical bonds. This dataset is designed to study node feature shifts, edge feature shifts, and structure shifts. The two covariate shift domains are scaffold graph type and molecular graph size. For the scaffold covariate shift, environments are partitioned based on the Bemis-Murcko scaffold (Bemis & Murcko, 1996), a two-dimensional structural base that does not determine a molecule’s ability to inhibit HIV replication. For the size covariate shift, the training distribution consists of molecular graphs ranging in size from $17$ to $222$ atoms. The validation set contains molecules with sizes between $15$ and $16$ atoms, while the test set includes molecules with sizes from $2$ to $14$ atoms.
  • GOOD-SST2: GOOD-SST2 is a real-world natural language sentimental analysis dataset from (Yuan et al., 2022), designed to investigate node feature shifts and structure shifts. Each graph is derived from a sentence, transformed into a grammar tree, where nodes represent words, and node features are corresponding word embeddings. The task is to predict the sentiment polarity of each sentence. Sentence length is chosen as the covariate shift environment, as sentence length should not inherently affect sentiment polarity. For the length covariate shift, the training distribution consists of grammar graphs with sizes ranging from $1$ to $7$ nodes, the validation distribution includes graphs with sizes from $8$ to $14$ nodes, and the test distribution contains graphs with sizes from $15$ to $56$ nodes.
A.4.2 Implementation settings
Diffusion models: Following (Jo et al., 2022), we preprocess each graph into two matrices: $\bm{X}\in\mathbb{R}^{n\times a}$ for node features, and $\bm{A}\in\mathbb{R}^{n\times n\times b}$ for adjacency and edge features. Here, $n$ represents the maximum number of nodes in a graph for the given dataset, while $a$ and $b$ denote the dimensions of node features and edge features, respectively. The graph structure, including edge features, is encoded in $\bm{A}$. For the GOOD-Motif dataset, $a$ corresponds to the node degree of a node. In GOOD-CMNIST, each node feature is the concatenation of its degree and color. In GOOD-SST2, the node feature is the word embedding. In the molecular dataset GOOD-HIV, $a$ represents possible atom types and $b$ denotes the types of bonds (e.g., single, double, triple). All molecules are converted to their kekulized form, with hydrogens removed using the RDKit library (Landrum et al., 2016). Additionally, we apply the valency correction proposed by (Zang & Wang, 2020) to post-process the generated molecules.
We train a graph transformer model (Dwivedi & Bresson, 2020; Vignac et al., 2022), $s_{\theta,t}$, to approximate the partial score functions $\nabla_{\bm{X}_{t}}\log p_{t}\left(\bm{X}_{t},\bm{A}t\right)$ and $\nabla{\bm{A}_{t}}\log p_{t}\left(\bm{X}_{t},\bm{A}_{t}\right)$ for the unlabelled graphs in the OOD training set and evaluate them on the OOD validation set. In line with (Jo et al., 2022), we use VP or VE SDEs to model the diffusion process for both node features and adjacency matrices. The specific details of the diffusion models are provided in Table 4.
@@ A.4.2 Implementation settings
We also train a graph transformer model, $\phi_{t}$, with the same architecture described in Table 4, to predict the class labels of the noisy graphs $G_{t}=\left(\bm{X}_{t},\bm{A}_{t}\right)$ at each time step $t$.
Table 4: Hyperparameters of diffusion models.

### v2

@@ 5.3. OOD Controlled Graph Generation
Visualization of OOD molecules We showcase the effectiveness of CODA by visualizing the OOD molecules generated by our approach in Table 4. As illustrated in Table 4, increasing $\lambda$ results in gradual modifications to the structures of the molecular graphs, while the properties of the molecules remain preserved. This controlled exploration allows CODA to produce chemically meaningful and diverse OOD molecules without disrupting their underlying validity. Notably, the changes remain smooth and interpretable across different $\lambda$ values, highlighting the tunable nature of our generation process.
Table 4. Visualizations of the augmented GOOD-Molhiv-scaffold graphs generated by CODA. The graphs $\tilde{G}$ with $\lambda=0.1$, $\lambda=0.2$, and $\lambda=0.3$ represent the OOD graphs generated by CODA under different values of $\lambda$.
6. Conclusion
We presented CODA, a score-based out-of-distribution graph generation framework for addressing covariate shift in graph learning. By coupling a conditional score with an exploration parameter, CODA explores low-density regions beyond the training support in a controlled manner while preserving label-determining stable patterns. Unlike edit-based augmentations, our approach jointly synthesizes structures, node features, and edge features and does not require explicit decomposition into stable and environmental components. Across synthetic, semi-artificial, and real benchmarks from GOOD, CODA consistently improves OOD generalization over invariant-learning and augmentation baselines. The exploration knob $\lambda$ yields monotonic increases in MMD while maintaining a high probability of retaining stable patterns, demonstrating both controllability and fidelity. Given that distribution shift is pervasive on the Web (e.g., evolving social, hyperlink, and interaction graphs), these results indicate practical value for robust web-scale modeling.
Overall, CODA provides a principled and controllable path to web-relevant graph augmentation, improving robustness to the distribution shifts that dominate modern web applications.
Appendix A Theory
A.1. Deriving the conditional reverse-time SDE
[…]
@@ A.1. Deriving the conditional reverse-time SDE
$\log p_{t}\left(G_{t},\mathbf{y}_{G}\mid\mathcal{E}_{\text{ood }}=\lambda\right)=$ $\log p_{t}\left(G_{t},\mathbf{y}_{G},\mathcal{E}_{\text{ood }}=\lambda\right)$
- $\log p_{t}\left(\mathcal{E}_{\text{ood }}=\lambda\right)$
Since $p_{t}\left(\mathcal{E}_{\text{ood }}=\lambda\right)$ is independent of $G_{t}$, $\nabla_{G_{t}}\log p_{t}\left(\mathcal{E}_{\text{ood }}=\lambda\right)=0$. Therefore,
[…]
@@
$\log p_{t}\left(G_{t}\mid\mathbf{y}_{G},\mathcal{E}_{\text{ood }}=\lambda\right)$ $=\log p_{t}\left(G_{t},\mathbf{y}_{G},\mathcal{E}_{\text{ood }}=\lambda\right)$
$-\log p_{t}\left(\mathbf{y}_{G},\mathcal{E}_{\text{ood }}=\lambda\right)$
Since $p_{t}\left(\mathbf{y}_{G},\mathcal{E}_{\text{ood }}=\lambda\right)$ is independent of $G_{t}$,
$\nabla_{G_{t}}\log p_{t}\left(\mathbf{y}_{G},\mathcal{E}_{\text{ood }}=\lambda\right)=0$
Therefore,
[…]
@@
[…]
Appendix B Metrics for measuring distributional differences
In this section, we provide detailed implementation steps for measuring the distributional differences between the augmented dataset and the training dataset. Following (thompson2022on), we first use an untrained random GIN, $h$, to extract graph embeddings from both the augmentation distribution and the training distribution. The maximum mean discrepancy (MMD) is then computed to quantify the dissimilarity between the graph embedding distributions:
$)]+\displaystyle\operatorname{MMD}^{2}(P\|\tilde{P})=\mathbb{E}_{g,\tilde{g}\sim P}[k(h(g),h(\tilde{g}$))]+
$)]\displaystyle\mathbb{E}_{g,\tilde{g}\sim\tilde{P}}[k(h(g),h(\tilde{g}))]-2\mathbb{E}_{g\sim P,\tilde{g}\sim\tilde{P}}[k(h(g),h(\tilde{g}$))]
where $k(\cdot,\cdot)$ is the RBF kernel proposed by (you2018graphrnn). As recommended by (thompson2022on), the MMD RBF scalar is also one of the most reliable metrics for measuring distributional differences:
@@ Appendix B Metrics for measuring distributional differences
Additionally, we employ the Earth Mover’s Distance (EMD) from (thompson2022on) to compute pairwise distances $d(\cdot,\cdot)$.
Appendix C Implementations of CODA.
Directly applying this framework to graphs is inadequate for addressing complex covariate shifts, especially those that simultaneously affect feature and structural distributions. To overcome this, we explicitly model the joint diffusion of node features and adjacency matrices, representing the perturbed graph as $\left\{G_{t}=\left(\boldsymbol{X}_{t},\boldsymbol{A}_{t}\right)\right\}_{t=0}^{T}$ using a set of SDEs for Equation 8:
(14) $\left\{\begin{aligned} \mathrm{d}\boldsymbol{X}_{t}=&{\left[\mathbf{f}_{1,t}\left(\boldsymbol{X}_{t}\right)-(1-\sqrt{\lambda})g_{1,t}^{2}\nabla_{\boldsymbol{X}_{t}}\log p_{t}\left(\boldsymbol{X}_{t},\boldsymbol{A}_{t}\right)\right.}\\ &\left.-(1-\sqrt{\lambda})g_{1,t}^{2}\nabla_{\boldsymbol{X}_{t}}\log p_{t}\left(\mathbf{y}_{G}\mid\boldsymbol{X}_{t},\boldsymbol{A}_{t}\right)\right]\mathrm{d}\bar{t}\\ &+g_{1,t}\mathrm{~d}\overline{\mathbf{w}}_{1}\\ \mathrm{~d}\boldsymbol{A}_{t}=&{\left[\mathbf{f}_{2,t}\left(\boldsymbol{A}_{t}\right)-(1-\sqrt{\lambda})g_{2,t}^{2}\nabla_{\boldsymbol{A}_{t}}\log p_{t}\left(\boldsymbol{X}_{t},\boldsymbol{A}_{t}\right)\right.}\\ &\left.-(1-\sqrt{\lambda})g_{2,t}^{2}\nabla_{\boldsymbol{A}_{t}}\log p_{t}\left(\mathbf{y}_{G}\mid\boldsymbol{X}_{t},\boldsymbol{A}_{t}\right)\right]\mathrm{d}\bar{t}\\ &+g_{2,t}\mathrm{~d}\overline{\mathbf{w}}_{2}.\end{aligned}\right$.
where $\mathbf{f}_{t}(\boldsymbol{X},\boldsymbol{A})=\left(\mathbf{f}_{1,t}(\boldsymbol{X}),\mathbf{f}_{2,t}(\boldsymbol{A})\right)$ and $g_{t}=\left(g_{1,t},g_{2,t}\right)$ representing the drift and diffusion coefficients, respectively. The reverse-time processes are captured by standard Wiener processes $\overline{\mathbf{w}}_{1}$ and $\overline{\mathbf{w}}_{2}$, with $\mathrm{d}\bar{t}$ indicating an infinitesimally small negative time step. We train one graph transformer (dwivedi2020generalization; vignac2022digress), denoted as $s_{\theta,t}=\left(s_{\theta_{1},t},s_{\theta_{2},t}\right)$ to closely estimate the partial score functions:
$s_{\theta,t}=\big(s_{\theta_{1},t},\,s_{\theta_{2},t}\big)\approx\Big(\nabla_{\boldsymbol{X}_{t}}\log p_{t}(\boldsymbol{X}_{t},\boldsymbol{A}_{t}),\;\nabla_{\boldsymbol{A}_{t}}\log p_{t}(\boldsymbol{X}_{t},\boldsymbol{A}_{t})\Big).$
Therefore, the denoising graph transformer is only trained with unlabelled graphs without $\lambda$ values and graph labels. The graph transformer takes perturbed node features $X_{t}$, adjacency matrices $A_{t}$ and normalized timestep as input. The timestep $t$ value is treated as a global graph feature, and an embedding layer is used to embed $t$.
@@
We also use a graph transformer model $\phi_{t}$ with the same architecture to predict the class label of the noisy graphs $G_{t}=\left(\boldsymbol{X}_{t},\boldsymbol{A}_{t}\right)$ at time step $t$. The target class $j$ probability $p_{t}\left(\mathbf{y}_{G}=j\mid\boldsymbol{X}_{t},\boldsymbol{A}_{t}\right)$ is then given by:
$p_{t}\left(\mathbf{y}_{G}=j\mid\boldsymbol{X}_{t},\boldsymbol{A}_{t}\right)=\frac{e^{\phi_{t}\left(\boldsymbol{X}_{t},\boldsymbol{A}_{t}\right)_{[j]}}}{\sum_{j=1}^{M}e^{\phi_{t}\left(\boldsymbol{X}_{t},\boldsymbol{A}_{t}\right)_{[j]}}}.$
Once both the score transformer and the classifier are trained, we use them to compute the conditional partial scores in Equation (8) during the sampling process. It is noteworthy that while our generation process leverages the representational power of graph transformers, subsequent evaluations for graph OOD classification rely on augmenting only simple GNN backbones with these generated samples (as detailed in Section 5.1).
[…]
where $\alpha_{t}=\left(\alpha_{1,t},\alpha_{2,t}\right)$, $r_{1}$ and $r_{2}$ are the weights for node features and adjacency matrices respectively, and $\|\cdot\|$ is the entry-wise matrix norm.
Intuitively, at the early stages of the reverse-time SDEs, the graphs are highly perturbed, resembling the prior noise distribution. Therefore, the classifier cannot accurately approximate the target class probability. Consequently, in the initial denoising steps, we focus more on guiding the reverse-time SDEs towards the low-density (i.e. OOD distribution) regions. As the reverse-time SDEs progressively denoise the graphs, we introduce guidance to direct the reverse-time SDEs towards regions exhibiting the desired stable patterns and the explored OOD environmental patterns.
Appendix D Experimental Details
D.1. Dataset Details
We utilize six datasets from the GOOD benchmark (gui2022good), including GOOD-Motif-base, GOOD-Motif-size, GOOD-CMNIST-color, GOOD-HIV-scaffold, GOOD-HIV-size, and GOOD-SST2-length. The GOOD benchmark (gui2022good) is the state-of-the-art framework for systematically evaluating graph OOD generalization. It carefully designs data environments to induce reliable and valid distribution shifts. The selected datasets span a diverse range of domains, covering covariate shifts in general graphs, image-transformed graphs, molecular graphs, and natural language sentiment analysis graphs. The dataset details are as follows:
  • GOOD-Motif: GOOD-Motif is a synthetic dataset from Spurious-Motif (wu2022discovering) specifically designed to investigate structure shifts. Each graph consists of an environmental base graph connected to a label-determining motif. The two primary covariate shift domains are the base graph type and graph size. For base covariate shift, the training distribution includes graphs with wheel, tree, and ladder base structures, while the validation set features star base graphs, and the test set contains path base graphs. For size covariate shift, the training distribution consists of graphs with sizes ranging from $6$ to $45$ nodes, the validation set contains graphs with sizes between $20$ and $75$ nodes, and the test set comprises graphs with sizes ranging from $68$ to $155$ nodes.
  • GOOD-CMNIST: GOOD-CMNIST is a semi-synthetic dataset designed to investigate node feature shifts. It consists of graphs transformed from MNIST handwritten digit images using superpixel techniques (monti2017geometric). Node color features are manually applied, making the color shift environment independent of the underlying structure. Specifically, for covariate shift, digits are colored using seven different colors. The training distribution includes digits colored with the first five colors, while the validation and test distributions contain digits with the remaining two colors, respectively.
  • GOOD-HIV: GOOD-HIV is a small-scale, real-world molecular dataset sourced from MoleculeNet (wu2018moleculenet). The nodes in these molecular graphs represent atoms, and the edges represent chemical bonds. This dataset is designed to study node feature shifts, edge feature shifts, and structure shifts. The two covariate shift domains are scaffold graph type and molecular graph size. For the scaffold covariate shift, environments are partitioned based on the Bemis-Murcko scaffold (bemis1996properties), a two-dimensional structural base that does not determine a molecule’s ability to inhibit HIV replication. For the size covariate shift, the training distribution consists of molecular graphs ranging in size from $17$ to $222$ atoms. The validation set contains molecules with sizes between $15$ and $16$ atoms, while the test set includes molecules with sizes from $2$ to $14$ atoms.
  • GOOD-SST2: GOOD-SST2 is a real-world natural language sentimental analysis dataset from (yuan2022explainability), designed to investigate node feature shifts and structure shifts. Each graph is derived from a sentence, transformed into a grammar tree, where nodes represent words, and node features are corresponding word embeddings. The task is to predict the sentiment polarity of each sentence. Sentence length is chosen as the covariate shift environment, as sentence length should not inherently affect sentiment polarity. For the length covariate shift, the training distribution consists of grammar graphs with sizes ranging from $1$ to $7$ nodes, the validation distribution includes graphs with sizes from $8$ to $14$ nodes, and the test distribution contains graphs with sizes from $15$ to $56$ nodes.
D.2. Implementation settings
Diffusion models: Following (jo2022score), we preprocess each graph into two matrices: $\boldsymbol{X}\in\mathbb{R}^{n\times a}$ for node features, and $\boldsymbol{A}\in\mathbb{R}^{n\times n\times b}$ for adjacency and edge features. Here, $n$ represents the maximum number of nodes in a graph for the given dataset, while $a$ and $b$ denote the dimensions of node features and edge features, respectively. The graph structure, including edge features, is encoded in $\boldsymbol{A}$. For the GOOD-Motif dataset, $a$ corresponds to the node degree of a node. In GOOD-CMNIST, each node feature is the concatenation of its degree and color. In GOOD-SST2, the node feature is the word embedding. In the molecular dataset GOOD-HIV, $a$ represents possible atom types and $b$ denotes the types of bonds (e.g., single, double, triple). All molecules are converted to their kekulized form, with hydrogens removed using the RDKit library (landrum2016rdkit). Additionally, we apply the valency correction proposed by (zang2020moflow) to post-process the generated molecules.
We train a graph transformer model (dwivedi2020generalization; vignac2022digress), $s_{\theta,t}$, to approximate the partial score functions for the unlabelled graphs in the OOD training set and evaluate them on the OOD validation set. In line with (jo2022score), we use VP or VE SDEs to model the diffusion process for both node features and adjacency matrices. The specific details of the diffusion models are provided in Table 5.
We also train a graph transformer model, $\phi_{t}$, with the same architecture described in Table 5, to predict the class labels of the noisy graphs $G_{t}=\left(\boldsymbol{X}_{t},\boldsymbol{A}_{t}\right)$ at each time step $t$.
Table 5. Hyperparameters of diffusion models.